OpenStudy (anonymous):

Can someone please explain and show me how to do these please. 1.) x^2 – 8x + 15 2.) a^2 – a – 20 3.)a^2 – 5a – 20 4.)a^2 + 12ab + 27b^2 5.)2a^2 + 30a + 100

OpenStudy (anonymous):

@Roya Could you help me?

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

all of 'em?

OpenStudy (anonymous):

dear i 'll . u need to find out x and a ?

OpenStudy (anonymous):

Yes please. Well, yeah. :/ It's factoring trinomials.

OpenStudy (anonymous):

Instructions say --> Factor each trinomial below. Please show your work

OpenStudy (anonymous):

x^2-8x+15=0 (x-3)(x-5)

OpenStudy (anonymous):

for ex for this one think about two numbers which their multiple product be 15 and their product of sum be -8

OpenStudy (anonymous):

we know that 3*5 is 15 and we check if the some of them is -8 or not ? 5+3=8 but with a little experience we guess -3+-5=-8

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

20=10*2 10*2=8 10-2=8 20=5*4 5+4=9 5-4=1 4-5=-1 so can u find out the factors ?

OpenStudy (anonymous):

10 and 5

OpenStudy (anonymous):

u mean (a-10)(a-5) ?

OpenStudy (anonymous):

Wait, which one are we on?

OpenStudy (anonymous):

The second one?

OpenStudy (anonymous):

second

OpenStudy (anonymous):

I'd say the factors would be -10 and -5.

OpenStudy (anonymous):

look at what i wrote before . 4-5=-1 4*-5=-20

OpenStudy (anonymous):

(a-1) (a-10)

OpenStudy (anonymous):

not -10 I meant a-1 * a-20

OpenStudy (anonymous):

a^2-11a+10 is it ur equation ?

OpenStudy (anonymous):

a^2-21a+20 is it ur equation

OpenStudy (anonymous):

So confusing. :/

OpenStudy (anonymous):

don't be confuse it's need just practice . look at what i done

OpenStudy (anonymous):

for second one as i wrote before the answer is : a-5*a+4

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

i think the third one is not correct but let's check forth one :

OpenStudy (anonymous):

The question for #2 is, a^2 – a – 20. So, I have to find the product of -a and the sum of -20, right?

OpenStudy (anonymous):

All those up top are questions, I have to factor all of them.

OpenStudy (anonymous):

u know (a+b)^2= a^2+2ab+b^2 can u guess what we should put in my equation instead b ? because u see that first term didn't change

OpenStudy (anonymous):

ok i solve num 3 as well but first the easier

OpenStudy (anonymous):

Let's do #2, again.

OpenStudy (anonymous):

a^2 – a – 20

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What would I do first?

OpenStudy (anonymous):

for factories as i remember we have some ways can u tell me the ways ?

OpenStudy (anonymous):

Find the product of one and the sum of another.

OpenStudy (anonymous):

plz tell me all the method

OpenStudy (anonymous):

I don't know..

OpenStudy (anonymous):

one is (x+b)(x+a)=x^2+(a+b)+ab that's what u mentioned .

OpenStudy (anonymous):

fortunately we can solve #2 with this : so look at 20 tell me product of which numbers make 20?

OpenStudy (anonymous):

10*2

OpenStudy (anonymous):

and 1- - 2 = 8

OpenStudy (anonymous):

10 - 2 = 8

OpenStudy (anonymous):

so 10 and 8?

OpenStudy (anonymous):

look at our formula coefficient of x is equal to sum of the numbers , am i correct ?

OpenStudy (anonymous):

Yse.

OpenStudy (anonymous):

ok now tell me is the sum of 10 and 2 equal to -1 ?

OpenStudy (anonymous):

10 + 2 = 12. 12 + (-13) = -1

OpenStudy (anonymous):

dear now what's ur number ? check them

OpenStudy (anonymous):

The sum of 10 and 2 is 12, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so the sum of 10 and 2 isn't -1.

OpenStudy (anonymous):

....

OpenStudy (anonymous):

dear don't be confuse . may i know what's ur grade ? high school ?

OpenStudy (anonymous):

9th.

OpenStudy (anonymous):

where r u from ?

OpenStudy (anonymous):

I just need to complete these 5 questions for my assignment or I get kicked out of my class. I'm from USA.

OpenStudy (anonymous):

dear i'm here till tomorrow morning and i'll help u . u r just 16 years old . never mind dear no one kick u out

OpenStudy (anonymous):

lets learn it again with each other ok ? ready ? fasten ur seat belt

OpenStudy (anonymous):

I'm 17. I've failed Algebra 3 times. I just don't understand it. I'd be graduated by now if it wasn't for Algebra.

OpenStudy (anonymous):

u r a genius . may i know ur name ?

OpenStudy (anonymous):

Possibly. What's yours?

OpenStudy (anonymous):

dear algebra is sweet if u touch it and guess it .

OpenStudy (anonymous):

my name is roya . i'm iranian and i'm 21 . i'm electrical engineering student and mu=y name means dream . any thing else ? i'm a girl

OpenStudy (anonymous):

Ha. I'm a boy, 17 the name is Dylan and I'm from Florida.

OpenStudy (anonymous):

ok are u ready ?

OpenStudy (anonymous):

My name is Careless because eeveryone calls me that and I don't care about anything but computers and programming and processing. Yes, I'm ready. Let's start with #2... again. please.

OpenStudy (anonymous):

dear look at # 2 . and compare it with our formula a^2-a-20

OpenStudy (anonymous):

our formula for this one is x^2+(a+b)x+ab

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

so we should find 2 numbers which their mlt = -20 and their sum = -1 ok ?

OpenStudy (anonymous):

mlt=multiple

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so tell me the numbers which their mlt = -20

OpenStudy (anonymous):

-10 * 2

OpenStudy (anonymous):

can u tell me their sum ?

OpenStudy (anonymous):

sum of...?

OpenStudy (anonymous):

-10 + 2?

OpenStudy (anonymous):

as u and me know their sum is 10-2 =8 ok ?

OpenStudy (anonymous):

or as u wrote -10+2=-8

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

none of them are equal to -1 . am i right ?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

so let's try another numbers . tell me two other number's which mlt=-20

OpenStudy (anonymous):

-5*4

OpenStudy (anonymous):

ok and tell me the sum

OpenStudy (anonymous):

-5 + 4 = -1///

OpenStudy (anonymous):

(x-5) (x+4) .... ^_^

OpenStudy (anonymous):

so u find two numbers which mlt=-20 and sum = -1 so put them in our formula

OpenStudy (anonymous):

such a genius boy . u should proud of urself .

OpenStudy (anonymous):

know multiple two terms to check ur answer .

OpenStudy (anonymous):

So, I got it right?

OpenStudy (anonymous):

ok i give u another ex and plz solve it for me .

OpenStudy (anonymous):

x^2+5x+6

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

mlt of 5 and sum of 6.. is....

OpenStudy (anonymous):

oh oh ! mlt=6 sum=5